# Two Lemmas on the Fiber-wise Holomorphicity in Complex Algebraic Geometry

### 12 Apr 2026 v1

@@ Abstract
Motivated by Hartogs’ theorem on separate holomorphicity, we establish two rigidity results in complex geometry.
First, we prove a lemma on regularity for weak solutions of algebraic differential equations, namely, we show that a fiber-wise distributional $L^{2}$ solution to an algebraic linear differential equation automatically upgrades to a global holomorphic solution on the total space, assuming the equation admits a compatible $L^{2}$ distributional solution along a transverse compact complex submanifold that maps birationally to the base.
Second, for a compact Kobayashi hyperbolic manifold fibered over $\mathbb{P}^{1}$, we prove that a continuous, fiber-wise holomorphic map of degree $1$ to a projective variety is a bi-holomorphic isomorphism, provided it is injective on a very ample hypersurface.
Keywords: Complex manifold, Hyperbolicity, Hartogs’ theorem, Birational geometry, Algebraic differential equation
@@ 1 Introduction and Background
In accordance with Hartogs’ theorem, we investigate the properties of fiber-wise holomorphic mappings.
We recall that in the theory of complex analysis, the Cauchy–Fantappiè integral formula is a fundamental tool for proving holomorphicity. However, transitioning from bounded domains to projective varieties presents a profound topological difference: compact complex manifolds possess no boundary. To replace the classical role of boundary integrals in forcing interior regularity, one must employ global algebraic geometry. In this paper, we present two related rigidity lemmas demonstrating how fiber-wise analytic behavior, when anchored by transverse algebraic intersections, such as rational sections and ample divisors, automatically elevates to global algebraic regularity.
2 A Lemma on Algebraic Differential Equations
In modern partial differential equation theory, it is standard to initially obtain solutions in a weak sense, such as $L^{2}$ distributions. The following result establishes a bridge, proving that if such weak fiber-wise solutions can be anchored by a compatible $L^{2}$ distributional solution on a transverse submanifold, their moduli are perfectly locked. This geometric anchoring forces the weak data to automatically patch together, bypassing the need for a priori global uniform estimates and elevating the solution to be globally holomorphic across the entire manifold.
Theorem 2.1.
Let $X,Y$ be smooth projective varieties over $\mathbb{C}$, and $f\colon X\rightarrow Y$ a surjective morphism with simply connected generic fibers. Let $M$ be a compact complex submanifold of $X$ such that $f|_{M}$ is birational. Let $\mathscr{E}\rightarrow X$ be a holomorphic vector bundle equipped with a flat connection $\nabla$, and $\sigma$ a holomorphic section of $\Omega^{1}_{X}\otimes\mathscr{E}$ satisfying $\nabla\sigma=0$. Suppose that there exist an $L^{2}$ section $\psi_{a}$ of $\mathscr{E}|_{f^{-1}(a)}$ for every regular value $a\in Y$ of $f$ and an $L^{2}$ section $\phi$ of $\mathscr{E}|_{M}$ solving the system
$\begin{cases}(i^{*}\nabla)\phi=\sigma|_{M},\\ (i_{a}^{*}\nabla)\psi_{a}=\sigma|_{f^{-1}(a)};\end{cases}$ (1)
in the sense of distributions, where $i\colon M\rightarrow X$ and $i_{a}\colon f^{-1}(a)\rightarrow X$ are the canonical inclusions. Then there exists a holomorphic solution $\Phi$ to the equation $\nabla\Phi=\sigma$ such that $\phi=\Phi|_{M}$.
@@ 2 A Lemma on Algebraic Differential Equations
By generic smoothness, there exists a Zariski dense open subset $U$ of $Y$, such that for every $a\in U$, the fiber $f^{-1}(a)$ is an irreducible smooth projective variety. Decomposing the connection as $\nabla=\nabla^{1,0}+\bar{\partial}$, the fiber-wise equations in system (1) imply $\bar{\partial}\phi=\bar{\partial}\psi_{a}=0$, as $\sigma$ is a $(1,0)$-form. By Weyl’s lemma on elliptic regularity applied to the $\bar{\partial}$-operator, the distributional solutions to system (1) are classical, and hence holomorphic.
Since the generic fibers of $f$ are simply connected, their fundamental groups vanish, making all flat vector bundles over $f^{-1}(a)$ trivializable. Any parallel section on $f^{-1}(a)$ is therefore uniquely determined by its evaluation at a single point without monodromy obstructions. Notice that, for a generic $a\in U$, the fiber $f^{-1}(a)$ intersects $M$ transversely. Since $f|_{M}$ is birational, there exists a Zariski dense open subset $V$ of $U$ over which $f|_{M}$ restricts to an isomorphism. By the vanishing of the obstruction class to solving the fiber-wise differential equation $(i_{a}^{*}\nabla)\psi=\sigma|_{f^{-1}(a)}$ and the holomorphic dependence on initial values, there exists a holomorphic local section $\Phi$ of $\mathscr{E}$ which solves the parametric family of complex PDEs
$\begin{cases}(i_{a}^{*}\nabla)\Phi|_{f^{-1}(a)}=\sigma|_{f^{-1}(a)},\\ \Phi|_{M}=\phi.\end{cases}$ (2)
For any $a\in V$, we have that $M$ intersects the fiber $f^{-1}(a)$ at exactly one point, rendering the initial value problem (2) well-posed. The local solution $\Phi$ thus extends to a global holomorphic section over $f^{-1}(V)$.
We now claim that the equality $\nabla\Phi=\sigma$ holds $f^{-1}(V)$. Consider the residual section $R:=-\sigma+\nabla\Phi$. Utilizing the flatness of the connection $\nabla$ alongside the closedness of $\sigma$, we deduce
$\nabla R=\nabla(\nabla\Phi-\sigma)=\nabla^{2}\Phi-\nabla\sigma=0.$
By the very construction, the restriction of $R$ to the fiber directions vanishes. Thus, for any vertical vector field $\xi$ tangent to the fibers, we have $R(\xi)=0$. Let $\eta$ be any local projectable horizontal vector field. Contracting $\nabla R$ with the bi-vector $\xi\otimes\eta$, we obtain
$(\nabla R)(\xi,\eta)=\nabla_{\xi}(R(\eta))-\nabla_{\eta}(R(\xi))-R([\xi,\eta])=0$
via Cartan’s formula for the exterior covariant derivative. Since $R$ annihilates vertical vectors, $\nabla_{\eta}(R(\xi))=0$. Because $\eta$ is projectable and $\xi$ is vertical, their Lie bracket $\xi,\eta]$ is purely vertical, implying $R([\xi,\eta])=0$ also vanishes, which reduces the equation to $\nabla_{\xi}(R(\eta))=0$, meaning the tensor contraction of $R$ with any horizontal vector field is a covariant constant along the fibers. Furthermore, when restricted to the transverse submanifold $M$, the tangent space splitting $T_{x}X=T_{x}M\oplus T_{x}(f^{-1}(a))$ forcing the tensor $R$ to be determined at $x$ by its pullbacks. Note that
$i^{*}R=(i^{*}\nabla)\Phi|_{M}-i^{*}\sigma=(i^{*}\nabla)\phi-\sigma|_{M}=0$
by the first equation in system (1). The transversality then yields that $R(\eta)$ vanishes at the intersection points of the fibers of $f$ with $M$. Since $R(\eta)$ is constant along the fibers and vanishes on $M$, it vanishes identically. Since $R(\xi)=R(\eta)=0$, we conclude that $R=0$, and hence the equation $\nabla\Phi=\sigma+R=\sigma$ holds on $f^{-1}(V)$. Because $\Phi$ satisfies $\nabla\Phi=\sigma$ globally on $f^{-1}(V)$, its local boundedness near the complement $X\setminus f^{-1}(V)$ follows from the holomorphic Poincaré lemma. Specifically, since $\nabla\sigma=0$, there exists a holomorphic local section s{s} satisfying $\nabla{s}=\sigma$ in a sufficiently small neighborhood of any point $x\in X\setminus f^{-1}(V)$. Consequently, we arrive at $\nabla(\Phi-{s})=0$ in the vicinity of $x$, which implies that $\Phi-{s}$ is a constant vector. The local boundedness of $\Phi$ is thus immediate. Consequently, the section $\Phi$ extends holomorphically across the proper analytic subset $X\setminus f^{-1}(V)$ to the entirety of $X$ by Riemann’s extension theorem, preserving the identity $\nabla\Phi=\sigma$ globally. ∎
Remark 2.2.
The assumption of an $L^{2}$ fiber-wise solution makes this lemma particularly adaptable to geometric analysis settings, allowing one to rigidly extend weak distributional limits directly without establishing global uniform continuity beforehand.
@@ 3 A Lemma on Kobayashi Hyperbolic Manifolds
Suppose now that $\dim(Y^{t})=n-2$. Take a generic point $y\in Y^{t}$. Then the fiber $C$ of $\phi^{t}$ over $y$ is a projective curve. Since $H\in\operatorname{div}(X)$ is a very ample divisor, $H^{t}:=X^{t}\cap H$ is very ample in $X^{t}$, making $H^{t}\cap C$ non-empty and $\phi(H^{t})=Y^{t}$. Because $\phi|_{H}$ is injective by assumption, the mapping degree of $\phi^{t}|_{H^{t}}$ onto its image is exactly 1. By choosing $y$ to be generic, we ensure that the restriction $\phi^{t}|_{H^{t}}$ is unramified at its unique preimage $x\in H^{t}$, rendering its differential an isomorphism at $x$. As a result, the tangent space $T_{x}H^{t}$ intersects trivially with the kernel of the differential of $\phi^{t}$ at $x$, which is precisely $T_{x}C$. This proves that $C$ intersects $H^{t}$ transversally, and hence their intersection product is $H^{t}\cdot C=1$. Because $H$ is a very ample divisor, intersecting a projective curve $C$ at a single point with a multiplicity of exactly 1 forces $C$ to be rational, that is, bi-holomorphic to $\mathbb{P}^{1}$. However, as $X$ is Kobayashi hyperbolic, there exists no non-constant holomorphic mapping from $\mathbb{P}^{1}$ into $X$ by Brody’s theorem on hyperbolicity. This is a contradiction. Therefore, we have in fact $\dim(Y^{t})=n-1$, so the regular mapping $\phi^{t}\colon X^{t}\rightarrow Y$ sends its domain of definition onto a projective variety $Y^{t}$ of equal dimension n1n-1. We claim that $\phi^{t}$ is actually a finite morphism. Consider the Stein factorization $\phi^{t}=\phi^{t}_{1}\circ\phi^{t}_{2}$ of $\phi^{t}$, where $\phi^{t}_{2}\colon X^{t}\rightarrow N$ is a surjective projective morphism with connected fibers onto a normal projective variety $N$, and $\phi^{t}_{1}\colon N\rightarrow Y^{t}$ is a finite morphism. Since $\phi^{t}_{1}$ is finite and surjective, we obtain
@@
Let $p_{1}\colon I\rightarrow X^{t_{1}}$ and $p_{2}\colon I\rightarrow X^{t_{2}}$ be the restrictions of the canonical projections onto each factor. Because $\phi^{t_{1}}$ and $\phi^{t_{2}}$ are finite morphisms, the projections $p_{1}$ and $p_{2}$ are also finite. Therefore, the pullbacks $p_{1}^{*}(H)$ and $p_{2}^{*}(H)$ are both ample divisors on the projective variety $I$. Since $\dim(I)\geq 2$, their intersection product $p_{1}^{*}(H)\cdot p_{2}^{*}(H)$ is strictly positive, ensuring the intersection of their support is non-empty. Take any $(h_{1},h_{2})\in p_{1}^{*}(H)\cap p_{2}^{*}(H)$. Then, by the very construction, we have $h_{1},h_{2}\in H$ such that $\phi(h_{1})=\phi(h_{2})$, which again contradicts the injectivity of $\phi|_{H}$. We conclude that the images $Y^{t}$ are mutually disjoint for all $t\in\Sigma$. Step 3: Because $Y$ is a smooth projective variety, and the collection $\{Y^{t}:t\in\Sigma\}$ consists of homologous effective divisors that are pairwise disjoint, the intersection product $Y^{t_{1}}\cdot Y^{t_{2}}$ vanishes for all distinct $t_{1},t_{2}\in\Sigma$. Therefore, a Zariski dense open subset of the compact complex manifold $Y$ is continuously foliated by the hypersurfaces $Y^{t}$ over $\Sigma$, yielding a continuous mapping
@@
Having established that $\phi\colon X\rightarrow Y$ is a proper holomorphic mapping of degree 1 between smooth projective varieties, it remains to show it is an isomorphism. But, if $\phi$ were not an isomorphism, as it is injective almost everywhere, it would be a birational map contracting an exceptional divisor to a lower-dimensional subvariety. Recall from birational geometry that the exceptional locus of a proper birational morphism between smooth projective varieties is uniruled, which once again contradicts the hyperbolicity of $X$.
Quod erat demonstrandum. ∎

### 16 Apr 2026 v2

# A Note on Two Useful Lemmas in Complex Algebraic Geometry

@@ Abstract
We establish two useful lemmas in complex geometry. First, we investigate the homological properties of differential operators intertwined by the Dolbeault Laplacian on compact Hermitian manifolds. By utilizing the mapping cone of Koszul complexes, we prove that the analytical obstructions to solving specific overdetermined PDE systems reduce entirely to the finite-dimensional Dolbeault cohomology. Second, for a compact Kobayashi hyperbolic manifold fibered over $\mathbb{P}^{1}$, we prove that a continuous, fiber-wise holomorphic map of degree $1$ to a projective variety is a bi-holomorphic isomorphism, provided it is injective on a very ample hypersurface.
Keywords: Dolbeault cohomology, Koszul complex, Complex manifold, Hyperbolicity
@@ 1 Introduction and Background
In accordance with the spirit of rigidity exemplified by Hartogs’ theorem, this paper investigates how rigid global structures in complex geometry can be enforced by partial algebraic or homological constraints. We divide our exploration into two distinct but philosophically related results.
In Section 2, we turn our attention to differential operators on compact Hermitian manifolds. In the study of integrable systems, operators intertwined by a Laplacian exhibit highly constrained solution spaces. We abstract this phenomenon using homological algebra. By establishing an acyclicity lemma for intertwining chain complexes, we demonstrate how the Dolbeault Laplacian acts as an invertible intertwining operator on the space of exact forms. This allows us to prove that the infinite-dimensional analytical obstructions to solving certain coupled PDE systems collapse entirely into the finite-dimensional Dolbeault cohomology.
In Section 3, we transition from the rigidity of differential equations to the rigidity of mappings. While Hartogs’ theorem relies on boundary integrals within bounded domains, transitioning to compact projective varieties requires global algebraic geometry. We present a rigidity lemma demonstrating how continuous, fiber-wise analytic mappings can be automatically elevated to global bi-holomorphic isomorphisms when anchored by a transverse algebraic intersection, specifically a very ample hypersurface, on a Kobayashi hyperbolic manifold.
2 A Lemma on Holomorphic Differential Equations
We begin our investigation by formulating a purely algebraic statement. By localizing the homological complexity to a specific submodule where an intertwining operator acts as an isomorphism, we equate the cohomology of the total space with that of the quotient.
Lemma 2.1.
Let $0\rightarrow A\rightarrow B\rightarrow C\rightarrow 0$ be a short exact sequence of $R$-modules, and let $f,g,h\in\operatorname{End}_{R}(B)$ satisfy $fg=gh$, so that the sequence $K(B)$ of $R$-modules given by
[…]
is a chain complex. Suppose that $h(A)\subseteq g(A)=A$ and $g|_{A}$ is a monomorphism. Then $0\rightarrow A\rightarrow A^{\oplus 2}\rightarrow A\rightarrow 0$ forms an acyclic subcomplex $K(A)$ of $K(B)$, such that the cohomology groups of $K(C):=K(B)/K(A)$ are isomorphic to that of $K(B)$.
@@ 2 A Lemma on Holomorphic Differential Equations
Since $fg=gh$, we deduce that $f(A)\subseteq A$. Thus, the endomorphisms $f,g,h$ restrict to the submodule $A$ and descend to the quotient $C=B/A$. We therefore obtain a short exact sequence
$0\longrightarrow K(A)\longrightarrow K(B)\longrightarrow K(C)\longrightarrow 0$
of complexes, which induces a long exact sequence
$\dots\longrightarrow H^{i}(K(A))\longrightarrow H^{i}(K(B))\longrightarrow H^{i}(K(C))\longrightarrow H^{i+1}(K(A))\longrightarrow\cdots$ (1)
of cohomology groups.
We claim that the subcomplex $K(A)$ is acyclic. To verify exactness at the first term, we take $a\in A$ satisfying g(a)=0-g(a)=0. Since $g|_{A}$ is a monomorphism, we obtain $a=0$. Next, to establish exactness at the middle term, we take $(a,b)\in A\oplus A$ such that $f(a)+g(b)=0$. Because $g(A)=A$ and $g|_{A}$ is injective, in fact $g|_{A}$ is an automorphism. Consequently, there exists a unique $c\in A$ such that g(c)=a-g(c)=a. Applying $g$ to $h(c)$ yields $g(h(c))=(gh)(c)$, which becomes f(g(c))=f(a)=f(a)f(g(c))=f(-a)=-f(a) by virtue of the intertwining relation $fg=gh$. Since $f(a)+g(b)=0$, we have g(b)=f(a)=g(h(c))g(b)=-f(a)=g(h(c)). Because $g|_{A}$ is injective, we conclude $b=h(c)$. Thus, it holds that (a,b)=(g(c),h(c))(a,b)=(-g(c),h(c)), proving exactness at $A\oplus A$. Finally, to prove exactness at the terminal term, we choose an arbitrary $c\in A$. Since $g|_{A}$ is surjective, there exists $b\in A$ such that $g(b)=c$, demonstrating surjectivity. The acyclicity of $K(A)$ is therefore established.
The exactness of sequence (1) then forces $H^{i}(K(B))=H^{i}(K(C))$ for all $i\in\mathbb{Z}$. ∎
We now translate this abstract homological machinery into the geometric language of differential equations on compact Hermitian manifolds. By choosing our short exact sequence to reflect the Hodge decomposition of differential forms, we let the Dolbeault Laplacian assume the role of the intertwining operator. Because the Laplacian acts invertibly on the orthogonal complement of the harmonic forms, Lemma 2.1 establishes a dimension reduction: the defects of operators are completely captured by the finite-dimensional Dolbeault cohomology.
Corollary 2.2.
Let $X$ be a compact Hermitian manifold, and $L,A$ linear differential operators on $E:=\Omega^{p,q}(X)\cap\operatorname{Ker}(\bar{\partial})$. Suppose that $A\Delta_{\bar{\partial}}=\Delta_{\bar{\partial}}L$ and $L(\operatorname{Im}(\bar{\partial})\cap E)\subseteq\operatorname{Im}(\bar{\partial})$. Then, the following properties hold:
  1. 1.
    a solution $\omega\in E$ to $L\omega=\Delta_{\bar{\partial}}\omega=0$ is precisely a harmonic form of which Dolbeault cohomology class lies in the kernel of the induced operator $L^{*}\colon H^{p,q}(X)\rightarrow H^{p,q}(X)$;
  2. 2.
    for any $\alpha,\beta\in E$ satisfying $A\alpha=\Delta_{\bar{\partial}}\beta$, there exists $\omega\in E$ simultaneously solving $\Delta_{\bar{\partial}}\omega=\alpha$ and $L\omega=\beta$ if and only if the cohomology class $\alpha]=0$ and $\beta]\in\operatorname{Im}(L^{*})$.
Proof.
Consider the short exact sequence of vector spaces
$0\longrightarrow\operatorname{Im}(\bar{\partial})\cap E\longrightarrow E\longrightarrow H^{p,q}(X)\longrightarrow 0.$
Since the Dolbeault Laplacian $\Delta_{\bar{\partial}}$ commutes with $\bar{\partial}$, it restricts to an endomorphism on $E$. By Hodge theory, we have $\Delta_{\bar{\partial}}(E)\subseteq\operatorname{Im}(\bar{\partial})\cap E$. Furthermore, as Green’s operator commutes with $\bar{\partial}$, the restriction of $\Delta_{\bar{\partial}}$ to the subspace $\operatorname{Im}(\bar{\partial})\cap E$ is an automorphism.
Moreover, for any $\alpha\in\operatorname{Im}(\bar{\partial})\cap E$, there exists $\omega\in\operatorname{Im}(\bar{\partial})\cap E$ such that $\alpha=\Delta_{\bar{\partial}}\omega$. Therefore, we obtain that
$A\alpha=A\Delta_{\bar{\partial}}\omega=\Delta_{\bar{\partial}}(L\omega).$
Since $L\omega\in E$, we have $\Delta_{\bar{\partial}}(L\omega)\in\Delta_{\bar{\partial}}(E)\subseteq\operatorname{Im}(\bar{\partial})\cap E$, which guarantees that the operator $A$ also preserves the subspace $\operatorname{Im}(\bar{\partial})\cap E$.
The hypotheses $A\Delta_{\bar{\partial}}=\Delta_{\bar{\partial}}L$ and $L(\operatorname{Im}(\bar{\partial})\cap E)\subseteq\operatorname{Im}(\bar{\partial})$ thus ensure that the operators under consideration satisfy all premises of Lemma 2.1. Using the notations in the proof of Lemma 2.1, we have the isomorphism $H^{i}(K(E))=H^{i}(K(H^{p,q}(X)))$ for all $i\in\mathbb{Z}$. Because $\Delta_{\bar{\partial}}(E)\subseteq\operatorname{Im}(\bar{\partial})\cap E$, the induced operator of $\Delta_{\bar{\partial}}$ on the quotient $H^{p,q}(X)$ is identically zero. The complex $K(H^{p,q}(X))$ thus simplifies to
[…]
Now, straightforward computation yields $H^{0}(K(H^{p,q}(X)))=\operatorname{Ker}(L^{*})$ and
$H^{1}(K(H^{p,q}(X)))=\operatorname{Ker}(A^{*})\oplus\operatorname{Coker}(L^{*}),$
which immediately translate to the two desired properties. ∎
@@ 3 A Lemma on Kobayashi Hyperbolic Manifolds
Suppose now that $\dim(Y^{t})=n-2$. Take a generic point $y\in Y^{t}$. Then the fiber $C$ of $\phi^{t}$ over $y$ is a projective curve. Since $H\in\operatorname{div}(X)$ is a very ample divisor, $H^{t}:=X^{t}\cap H$ is very ample in $X^{t}$, making $H^{t}\cap C$ non-empty and $\phi(H^{t})=Y^{t}$. Because $\phi|_{H}$ is injective by assumption, the mapping degree of $\phi^{t}|_{H^{t}}$ onto its image is exactly 1. By choosing $y$ to be generic, we ensure that the restriction $\phi^{t}|_{H^{t}}$ is unramified at its unique preimage $x\in H^{t}$, rendering its differential an isomorphism at $x$. As a result, the tangent space $T_{x}H^{t}$ intersects trivially with the kernel of the differential of $\phi^{t}$ at $x$, which is precisely $T_{x}C$. This proves that $C$ intersects $H^{t}$ transversally, and hence their intersection product is $H^{t}\cdot C=1$. Because $H$ is a very ample divisor, intersecting a projective curve $C$ at a single point with a multiplicity of exactly 1 forces $C$ to be rational, that is, bi-holomorphic to $\mathbb{P}^{1}$. However, as $X$ is Kobayashi hyperbolic, there exists no non-constant holomorphic mapping from $\mathbb{P}^{1}$ into $X$ by Brody’s theorem on hyperbolicity. This is a contradiction. Therefore, we have in fact $\dim(Y^{t})=n-1$, so the regular mapping $\phi^{t}\colon X^{t}\rightarrow Y$ sends its domain of definition onto a projective variety $Y^{t}$ of equal dimension n1n-1.
We claim that $\phi^{t}$ is actually a finite morphism. Consider the Stein factorization $\phi^{t}=\phi^{t}_{1}\circ\phi^{t}_{2}$ of $\phi^{t}$, where $\phi^{t}_{2}\colon X^{t}\rightarrow N$ is a surjective projective morphism with connected fibers onto a normal projective variety $N$, and $\phi^{t}_{1}\colon N\rightarrow Y^{t}$ is a finite morphism. Since $\phi^{t}_{1}$ is finite and surjective, we obtain
@@
Let $p_{1}\colon I\rightarrow X^{t_{1}}$ and $p_{2}\colon I\rightarrow X^{t_{2}}$ be the restrictions of the canonical projections onto each factor. Because $\phi^{t_{1}}$ and $\phi^{t_{2}}$ are finite morphisms, the projections $p_{1}$ and $p_{2}$ are also finite. Therefore, the pullbacks $p_{1}^{*}(H)$ and $p_{2}^{*}(H)$ are both ample divisors on the projective variety $I$. Since $\dim(I)\geq 2$, their intersection product $p_{1}^{*}(H)\cdot p_{2}^{*}(H)$ is strictly positive, ensuring the intersection of their support is non-empty. Take any $(h_{1},h_{2})\in p_{1}^{*}(H)\cap p_{2}^{*}(H)$. Then, by the very construction, we have $h_{1},h_{2}\in H$ such that $\phi(h_{1})=\phi(h_{2})$, which again contradicts the injectivity of $\phi|_{H}$. We conclude that the images $Y^{t}$ are mutually disjoint for all $t\in\Sigma$.
Step 3: Because $Y$ is a smooth projective variety, and the collection $\{Y^{t}:t\in\Sigma\}$ consists of homologous effective divisors that are pairwise disjoint, the intersection product $Y^{t_{1}}\cdot Y^{t_{2}}$ vanishes for all distinct $t_{1},t_{2}\in\Sigma$. Therefore, a Zariski dense open subset of the compact complex manifold $Y$ is continuously foliated by the hypersurfaces $Y^{t}$ over $\Sigma$, yielding a continuous mapping
@@
Having established that $\phi\colon X\rightarrow Y$ is a proper holomorphic mapping of degree 1 between smooth projective varieties, it remains to show it is an isomorphism. But, if $\phi$ were not an isomorphism, as it is injective almost everywhere, it would be a birational map contracting an exceptional divisor to a lower-dimensional subvariety. Recall from birational geometry that the exceptional locus of a proper birational morphism between smooth projective varieties is uniruled, which once again contradicts the hyperbolicity of $X$. ∎